OpenStudy (anonymous):

how is it that a lot of people come up with 2? is it because they dropped the brackets after adding the 9+3?

OpenStudy (anonymous):

order of operations

OpenStudy (anonymous):

how is it that the answer is 2??? are you talking about the same previous question?

OpenStudy (anonymous):

Its because the multiply the 2 by the 12 and get 48/24 but thats not how its works. For you to multiply the 2 by the 12 it would have to look like 49/(2*(9+3))

OpenStudy (anonymous):

48/2(9+3)????

OpenStudy (anonymous):

thats not how it looks....it looks like this 48÷2 (9+3) = ?

OpenStudy (anonymous):

Then its very ambiguous and i would argue with your teacher.

OpenStudy (anonymous):

yeah, it all depends how u look at it. Is this a math problem from a book or one that the teacher gave out?

OpenStudy (anonymous):

If i were to do this problem I would see it as (48/2)*11

OpenStudy (anonymous):

That would be 264

OpenStudy (anonymous):

are your classmates getting two????

OpenStudy (anonymous):

How do you get 11....... 9+3 is 12, i think we can agree those brackets are done first.

OpenStudy (anonymous):

cause if they did it this way then: 48/(2*11) will give you 48/22

OpenStudy (anonymous):

wow, my bad. I need to check my eyes........lol

OpenStudy (anonymous):

yeah so it would be (48/2)12 which would be 288

OpenStudy (anonymous):

Lol, i feel you. My calculus studying is boring me so i discovered this place and yet ive made some simple errors. Grrrrrr being tired, having bad eyes and such.

OpenStudy (anonymous):

yeah, I stayed up late last night studying, so i'm not my self right now, lol.

OpenStudy (anonymous):

did you find this site through paul's online notes???

OpenStudy (anonymous):

thank you one that the teacher gave out yes it would be 9+3 =12 and 48 /2=24 and 24/12 =2

OpenStudy (anonymous):

Yeah, Studying some triple integrals atm. Great study guide when i barely payed attention all year.

OpenStudy (anonymous):

that is how all of my classmates are seeing it

OpenStudy (anonymous):

you know what I remember one of my teacher doing one of these problems where a bunch of us go it wrong because we had to stick to the order of operations.

OpenStudy (anonymous):

you know, PEMDAS.

OpenStudy (anonymous):

@OP You should understand that your teachers question was ambiguous. When you get to higher up math, put stuff in a line as such 48/2*(9+3) means that only 48 is divided by the 2. and not by the 12 as well. You would need brackets surrounding the 2*(9+3) for it to be divided by that as well ie 48/(2*(9+3)). but this is low level math where many things are taught so that it is easy to teach. Later on you will relearn ALOT.

OpenStudy (anonymous):

yes...please excuse my dear aunt sally

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

I would agree with finally31, things are more specific later one. You will definitely know what your supposed to do.

OpenStudy (anonymous):

Just ask your teacher for more clarification on this.

myininaya (myininaya):

I'm confused on why everyone is confused about the order of operations. Do the order of operations. so we have 48 divided by 2(9+3)=48 divided by 2(12)=24(12)=288. Just simply follow the order of operations. It isn't ambiguous. After you perform the operation inside the grouping symbols, you do the division and multiplication as it occurs from left to right. So which ever occurs first do that first.